how do you solve y=4x+2 and y=3x+6 step by step?
Answer by ewatrrr(24785)

 
Hi,
y=4x+2 and
y=3x+6
therefore:
4x+2 = 3x+6
Solving for x
x = 4 and y = 18 (3*4+6)
Ordered pair (4,18) the solution for this system of equations

CHECKING our Answer***
y=4x+2
18 = 4*4 + 2